Chap 6 #3
Please complete the tables. Thank you!!!

3 lnormal variable Z (You may find it useful to reference the z table nd your answers to 4 decimal places.) 5 points a. P(Z> 1 05) b P(Z s -1.7) c POs2 1.46) eBook d P(-0.59 s Z s 2 86) Hint

And


Find the following z values for the standard normal variable Z (You may find it useful to reference the ztable. Negative values should be indicated by a minus sign. Round your answers to 2 decimal places) b P(Z> z) 0.742
Answer #1

3. P

robability values :

a. 0.1469
b. 0.0446
c. 0.4279
d. 0.7203

z values:

a.1.23
b. -0.65
c. 2.17
d. 1.04

ol. 0. 921855 -o.5 0-427 9 0 . 99 738ー0-27ク595 0·ワ203 =

b. pf久)%)=0.742 1.Pİ2sz) = o.TAa 5 8Plzsz)= 0.485+0.5 plz z) =0.985 ng table vala, ae get r(İzle 2.17009):0.985 2.17 z=1.0千